OpenStudy (anonymous):

TRUE or FALSE?! :) The law of cosines can be applied to right and non-right triangles. ***MY answer: TRUE! is that right? :)

OpenStudy (hunus):

Yes

OpenStudy (anonymous):

yay! Thanks :)

OpenStudy (hunus):

The right triangle is a special case of the law of cosines \[\Large c^2 = a^2 + b^2 - 2AB\cos(\theta)\] \[\Large c^2 = a^2 + b^2 - 2AB\cos(\frac{\pi}{2})\] \[\Large c^2 = a^2 + b^2\]